Circular No. 5775

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 J. A. Phillips and S. R. Kulkarni, California Institute of Technology,

report: “Using the four-element millimeter interferometer at the Owens Valley Radio Observatory (OVRO), we have continued monitoring the flux density of SN 1993J at a frequency of 99.4 GHz (± 3 mJy): Apr. 11.24-11.52 UT, 18 mJy; 14.23-14.53, 17; 21.22-21.52, 20. The fluxes listed for Apr. 11 and 14 supersede the values we reported on IAUC 5763; the previous fluxes were based on continuum synthesis maps with an effective bandwidth of 250 MHz. The new values cited here were obtained from a wideband digital correlator with an effective bandwidth of 1 GHz, and the resulting uncertainties are a factor of two smaller. Our flux scale was established by observations of Uranus and Neptune. The respective fluxes we measured during the supernova observations for three other compact radio sources are as follows: 3C 273, 22.1, 25.4, 27.3 Jy; 0954+658, 1.9, 1.9, 2.2 Jy; 3C 345, –, 5.1, 5.9 Jy. This will facilitate comparisons of our flux scale with that of other millimeter telescopes currently engaged in monitoring the supernova. The continuum source 0954+658 is only 24′ from the position of the supernova and was used as a phase calibrator for the interferometric observations. We have also used the OVRO millimeter array to observe the supernova at 221.9 GHz. Our first attempt was on Apr. 20, and we did not detect any source at the location of the radio supernova above a 3σ threshold of 48 mJy.”

B. Skiff, Lowell Observatory, provides the following Johnson V magnitudes and Strömgren b–y colors (± 0.02) obtained on Apr. 26.1 UT for several additional GSC stars 4383.n near the supernova (cf. IAUC 5742): 0150, V = 10.45, b–y = +0.79; 0384, 11.01, +0.33 (see note below); 0565, 11.42, +0.36; 0698, 10.82, +0.66. The GSC V magnitudes of these stars are 9.87, 10.85, 12.33, and 10.37, respectively. On Apr. 26.15, the supernova was at V = 11.35, b–y = +0.88.

D. Hanzl, N. Copernicus Planetarium, Brno, communicates that GSC 4383.0384 appears to be variable, first noted visually on Apr. 16 by K. Hornoch at Lelekovice and J. Kysely at Vlasim, and confirmed photoelectrically by Hanzl as having a maximum at V = 10.98 ± 0.05, amplitude ∼ 0.35 mag, and period of probably < 1 day. Hanzl also provides the following additional Johnson photometry for SN 1993J: Apr. 20.89 UT, V = 10.88 ± 0.02, B–V = +0.74 ± 0.02; 23.92, 11.09 ± 0.05, +1.00 ± 0.06; 25.02, 11.22 ± 0.04, +1.05 ± 0.04.

1993 April 26 Daniel W. E. Green